OpenStudy (honey522):

Write the equation of the line in point-slope form that passes through (5, -1) and has a slope of 2/3. A.) y-1=2/3 (x+5) B.)y-1=2/3 (x-5 C.)y+1=2/3 (x-5) D.)y+1=2/3(x+5

OpenStudy (dayakar):

equation of a line whose slope (m) = 2/3 and passing through the point x1 ,y1) = (5 ,-1) is y-y1 = m(x-x1)

OpenStudy (ijlal):